OpenStudy (anonymous):

What is the limit as x->0 of sin5x/sin4x?

OpenStudy (anonymous):

5/4

OpenStudy (anonymous):

Wait it isnt 4/5? thats what i was thinking

OpenStudy (anonymous):

nope its 5/4

OpenStudy (anonymous):

Alright. any idea why?I try it and get 4/5?

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sin(Ax)}{\sin(Bx)}=\lim_{x \rightarrow 0}\frac{A}{B} \cdot \frac{Bx}{Ax} \cdot \frac{\sin(Ax)}{\sin(Bx)}\] \[=\lim_{x \rightarrow 0}\frac{A}{B}\frac{\sin(Ax)}{Ax} \cdot \frac{Bx}{\sin(Bx)}=\frac{A}{B} \cdot 1 \cdot 1 \]

OpenStudy (anonymous):

ah i think i see it. Thanks myininaya for the explaination

myininaya (myininaya):

np so this case i provided you is a general one

myininaya (myininaya):

you understand it?

OpenStudy (anonymous):

yes I think so. It would be 5/4 because its just 5/4 multipled by 1 twice?

OpenStudy (anonymous):

MAybe I dont get it. Not sure how you get to the first part then the 2nd.

myininaya (myininaya):

yes A=5 B=4

myininaya (myininaya):

i just multiplied by 1

myininaya (myininaya):

\[1=\frac{ABx}{ABx}\]

myininaya (myininaya):

then i rearrange so i could write sin(Ax)/Ax and Bx/sin(Bx)

myininaya (myininaya):

and then i had A/B left

OpenStudy (anonymous):

why did you multiple by A/B and Bx/Ax?

myininaya (myininaya):

i know sin(Ax)/Ax->1 as x->0 i know Bx/sin(Bx)->1 as x->0 so i was thinking how do i get this Bx/Ax i have to also multiply by A/B

OpenStudy (anonymous):

Yeah I dont see it at all really..Where did you get the first part from and how did that equal the second part?I dont see how it just suddenly equals sin(ax)/ax

myininaya (myininaya):

\[\frac{ABx}{ABx} \cdot \frac{\sin(Ax)}{\sin(Bx)}=\frac{A}{B} \cdot \frac{Bx}{Ax} \cdot \frac{\sin(Ax)}{\sin(Bx)}\] \[=\frac{A}{B} \cdot \frac{\sin(Ax)}{Ax} \cdot \frac{Bx}{\sin(Bx)}\]

myininaya (myininaya):

so you don't understand this?

myininaya (myininaya):

i just moved stuff around

myininaya (myininaya):

remember multiplication is commutative

OpenStudy (anonymous):

Yeah I'm not seeing it... I dont get how Sin5x/sin4x = (5/4)sinx5x/5x)(4x/sin4x)

OpenStudy (anonymous):

It just looks like there is way too much there

myininaya (myininaya):

\[\frac{\sin(5x)}{\sin(4x)}=\frac{5(4)x}{5(4)x} \cdot \frac{\sin(5x)}{\sin(4x)}=\frac{\sin(5x)}{5x} \cdot \frac{4x}{\sin(4x)} \cdot \frac{5}{4}=1(1)(5/4)=5/4\]

myininaya (myininaya):

and thats of course if x->0 then L=5/4

myininaya (myininaya):

i don't understand what part you aren't understanding

OpenStudy (anonymous):

Oh I guess

myininaya (myininaya):

its just rearranging

myininaya (myininaya):

a(b)=b(a)

myininaya (myininaya):

you know this right?

myininaya (myininaya):

example: 2(4)=4(2)

myininaya (myininaya):

what does sin(x)/x go to as x goes to 0?

OpenStudy (anonymous):

Yes I guess it now. Just havent really done rearranging with fractions like that. Wasnt sure if we could

OpenStudy (anonymous):

Yes I guess it now. Just havent really done rearranging with fractions like that. Wasnt sure if we could

myininaya (myininaya):

\[\lim_{u(x) \rightarrow 0}\frac{\sin(u(x))}{u(x)}=1\]

myininaya (myininaya):

 right?